OpenStudy (anonymous):

Need to know how to solve, (432w+2592)/(w*w+12w+36) = ((432w)/(w*w))-1 answer is w= 54,48

OpenStudy (anonymous):

\[{(432w+2592)\over(w^2+12w+36)} ~~~~= {{{(432w)}\over {(w^2)}}~~-1} \]

OpenStudy (anonymous):

simplify the trinomial

OpenStudy (anonymous):

(432w+2592)/((w+6)^2)=(432-w^2)/(w^20

OpenStudy (anonymous):

okay multiply EACH TERM by \[\frac{ \left(\begin{matrix}\\(w ^{2})(x+6)^{2} \end{matrix}\right) }{ 1}\]

OpenStudy (anonymous):

but simplify by crossing out what yu can befor multiplying.

OpenStudy (anonymous):

I got this but pretty sure it's wrong: 2592w^4+31104w^5+93312w^2 = -w^6 - 12w^4-36w^4

OpenStudy (anonymous):

\[(w ^{2})(432w+2592)=(w+6)^{2}(432w)-(w ^{2})(w+6)^{2}\] simplify

OpenStudy (anonymous):

unless thats what yu got already

OpenStudy (anonymous):

432w^3+2592w^2 = 432w^3+432w^2+15552w-w^4+12w^3+36w^2

OpenStudy (anonymous):

ok the w+6 needs to be put back into the trinomial first

OpenStudy (anonymous):

so do (w^2)(w^2+12w+36) times each term?

OpenStudy (anonymous):

no thats already been done. im talking about this part 432w^3+432w^2+15552w-w^4+12w^3+36w^2 did yu multiply the 432w by the trinomial?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay then simplify all.

OpenStudy (anonymous):

set it equal to zero or just simplify both sides?

OpenStudy (anonymous):

get the w one one side. or yeah set to 0

OpenStudy (anonymous):

-w^4-12w^3-2196w^2+15552w=0

OpenStudy (anonymous):

can yu simplify more?

OpenStudy (anonymous):

not that I know of

OpenStudy (anonymous):

a 'w'maybe?

OpenStudy (anonymous):

I'm not sure how to simplify it any further

OpenStudy (anonymous):

factor out the w.

OpenStudy (anonymous):

so yu get w(whatever you get in the parenthesis)=0

OpenStudy (anonymous):

w(-1^4-12^3-2196^2+15552)=0

OpenStudy (anonymous):

no. the only thing that change is the exponents. they all go down by one.

OpenStudy (anonymous):

w(-1^3-12^2-2196+15552)=0 is that what u mean?

OpenStudy (anonymous):

no the w stay where they are.

OpenStudy (anonymous):

so the first term is -w^3

OpenStudy (anonymous):

-w^3-12^2-2196w+15552=0

OpenStudy (anonymous):

all w's stay. even on the ^2

OpenStudy (anonymous):

-w^3-12w^2-2196w+15552=0 That?

OpenStudy (anonymous):

dont forget yur w( ) but keep factoring the polynomial.

OpenStudy (anonymous):

w(-w^3-12w^2-2196w+15552)=0 But how can I factor more than that?

OpenStudy (anonymous):

we can try factor by grouping right?

OpenStudy (anonymous):

I don't know that method unless I call it something else..

OpenStudy (anonymous):

its algebra 1 factoring.

OpenStudy (anonymous):

so divide the exponents by w?

OpenStudy (anonymous):

wait a min....

OpenStudy (anonymous):

okay yeah w(-w^3-12w^2)(-2196w+15552)=0 and factor out the parenthesises

OpenStudy (anonymous):

Do I distribute the w first.

OpenStudy (anonymous):

no. take the set and factor it

OpenStudy (anonymous):

w(-w-12)(-2196+15552)=0 eh?

OpenStudy (anonymous):

okay lets just work on the first set of the parenthesis. (-w^3-12w^2)

OpenStudy (anonymous):

Ok, I'm not sure what to do if what I did before isn't right

OpenStudy (anonymous):

yu should get -w^2(w+12)

OpenStudy (anonymous):

ah ok, that makes sense

OpenStudy (anonymous):

can yu do the second term?

OpenStudy (anonymous):

36(-61w+432) ?

OpenStudy (anonymous):

right. now the -w^2 will be to the -36 so yu yuterms and -w^2-36 and(-61w+432) (w+12)

OpenStudy (anonymous):

now do I combine them?

OpenStudy (anonymous):

pretty much.but seperate them all by parenthesis.

OpenStudy (anonymous):

-w^2-60w+408=0 ?

OpenStudy (anonymous):

no. w( -w^2-36)(-61w+432)(w+12)=0

OpenStudy (anonymous):

got it?